# Discriminating QCD Compton and Quark–Antiquark Annihilation Processes in $\gamma$+Jets Using Interpretable Machine Learning

### v1

@@ II.1 Event generation and Jet reconstruction
Monte Carlo $pp$ events at $\sqrt{s}$ = 13 TeV are generated using the PYTHIA8-Detroit tune [8] with $\hat{p}_{\rm T}>$ 30 $\rm GeV/{\it c}$. Two classes of events are produced: inclusive dijet events and $\gamma+{\rm jet}$ events. Jet flavor labels are assigned based on the flavor of the highest-$p_{T}$ outgoing parton matched to the reconstructed jet from different processes. On the other hand, the $\gamma+{\rm jet}$ samples are used to study the discrimination between QCD Compton and quark–antiquark ($q-\bar{q}$) annihilation processes based on the flavor of the recoiling jet. Parton showering, hadronization, and underlying-event effects are modeled using the default PYTHIA8-Detroit tune. All results presented in this work are based on simulations generated with PYTHIA8-Detroit tune; no attempt is made to assess tune-dependent variations or other MC event generator, which are left for future studies.
Jets are reconstructed from final-state particles using anti-$k_{\rm T}$ jet-clustering algorithm provided in FastJet package [11]. Jets are reconstructed with a resolution parameter $R=0.5$, which provides a reasonable balance between retaining jet radiation and minimizing sensitivity to soft contamination; a systematic study of the jet-radius dependence is left for future work. The same jet reconstruction procedure and parameters are applied consistently to both the dijet samples used for training and the $\gamma+{\rm jet}$ samples used for process discrimination. This ensures that differences in classifier performance arise from the underlying jet structure rather than from reconstruction-related effects. For a given jet flavor, the jet substructure observables show consistent distributions between dijet and $\gamma+$jet events, indicating that the features learned from dijet samples are transferable to $\gamma+$jet recoil jets (Appendix A).
@@ II.2 Jet Observables
Figure 1 shows the distributions of the key input observables— $p_{\rm T,jet}$, $M_{\rm jet}$, $Q^{\rm ch}_{\kappa}$, $N_{\rm const}$, and girthgirth— used to train the classifier to distinguish quark- and gluon-initiated jets in dijet events. Jets are selected with $p_{\rm T,jet}$¿ 30 $\rm GeV/{\it c}$, and the jet substructure observables $M_{\rm jet}$, $Q^{\rm ch}_{\kappa}$, $N_{\rm const}$, and girthgirth are computed for these jets. These observables probe complementary aspects of the jet radiation pattern, like the overall QCD radiation, and radial energy distribution. While the distributions exhibit substantial overlap, systematic differences between quark- and gluon-initiated jets are observed, providing discrimination power on a statistical basis. The overlap between the quark and gluon distributions reflects the intrinsic limitations of quark–gluon discrimination arising from QCD radiation,motivating the use of multivariate techniques to exploit correlations among these observables.
@@ II.3.1 Multilayer Perceptron (MLP)
Figure 4: Using the MLP model: $p_{\rm T,jet}$, girthgirth, $M_{\rm jet}$, $N_{\rm const}$, and $Q^{\rm ch}_{\kappa}$distributions are shown for $\gamma+{\rm jet}$ with 50 << $p_{\rm T,jet}$<< 100 $\rm GeV/{\it c}$. Here, the $y$-axis is normalized by the total number of jets (NN). The blue hashed region and blue marker represent the true quark and predicted quark-like jet (from QCD Compton process); the same for red color represents gluon jet (from the annihilation process).
@@ II.3.2 Boosted Decision Tree (BDT)
In the BDT classifier, gradient boosting is employed with an ensemble of 1000 decision trees, each with a maximum depth of five. To improve stability and reduce overtraining, a learning rate (shrinkage) of 0.05 is used together with bagging, where each tree is trained on a random 50% subset of the training sample. Node splits are optimized using the Gini index as the separation criterion, and cost–complexity pruning is applied to further regularize the trees. These settings provide a balance between classification performance and robustness; the ROC is discussed in the result III.
@@ III Results and discussion
The close agreement between the BDT and MLP results indicates that the dominant discriminating information is already captured by the selected set of jet observables, with the observed performance saturation driven by physics rather than algorithmic limitations. In future, it would be important to incorporate the higher-dimensional jet substructure information to correlate QCD effects and to implement the advanced machine learning approach. These approaches are particularly well suited to precision jet studies at the LHC and the future EIC, where controlled kinematics and reduced backgrounds enable detailed exploration of correlated QCD dynamics.

### v2

@@ II.1 Event generation and Jet reconstruction
Monte Carlo $pp$ events at $\sqrt{s}$ = 13 TeV are generated using the PYTHIA8-Detroit tune [8] with $\hat{p}_{\rm T}>$ 30 $\rm GeV/{\it c}$. Two classes of events are produced: inclusive dijet events and $\gamma+{\rm jet}$ events.
Jets are reconstructed from final-state particles using anti-$k_{\rm T}$ jet-clustering algorithm provided in FastJet package [12]. Jet resolution parameter $R=0.5$ is used, which provides a reasonable balance between retaining jet radiation and minimizing sensitivity to soft contamination; a systematic study of the jet-radius dependence is left for future work. Jet flavor labels are assigned based on the flavor of the highest-$p_{T}$ outgoing parton matched with to the reconstructed jet from different processes. Here the matching criterion is $\Delta R<$0.25; the $\Delta R$ is the separation between the hard parton and jet axis in $\eta-\phi$ plane. The $\gamma+{\rm jet}$ samples are used to study the discrimination between QCD Compton and quark–antiquark ($q-\bar{q}$) annihilation processes based on the flavor of the recoiling jet. Parton showering, hadronization, and underlying-event effects are modeled using the default PYTHIA8-Detroit tune. All results presented in this work are based on simulations generated with PYTHIA8-Detroit tune; no attempt is made to assess tune-dependent variations or other MC event generator, which are left for future studies.
The same jet reconstruction procedure and parameters are applied consistently to both the dijet samples used for training and the $\gamma+{\rm jet}$ samples used for process discrimination. This ensures that differences in classifier performance arise from the underlying jet structure rather than from reconstruction-related effects. For a given jet flavor, the jet substructure observables show consistent distributions between dijet and $\gamma+$jet events, indicating that the features learned from dijet samples are transferable to $\gamma+$jet recoil jets (Appendix A).
@@ II.2 Jet Observables
Figure 1 shows the distributions of the key input observables— $p_{\rm T,jet}$, $M_{\rm jet}$, $Q^{\rm ch}_{\kappa}$, $N_{\rm const}$, and girthgirth— used to train the classifier to distinguish quark- and gluon-initiated jets in dijet events. Jets are selected with $p_{\rm T,jet}$>> 30 $\rm GeV/{\it c}$, and the jet substructure observables $M_{\rm jet}$, $Q^{\rm ch}_{\kappa}$, $N_{\rm const}$, and girthgirth are computed for these jets. These observables probe complementary aspects of the jet radiation pattern, like the overall QCD radiation, and radial energy distribution. While the distributions exhibit substantial overlap, systematic differences between quark- and gluon-initiated jets are observed, providing discrimination power on a statistical basis. The overlap between the quark and gluon distributions reflects the intrinsic limitations of quark–gluon discrimination arising from QCD radiation,motivating the use of multivariate techniques to exploit correlations among these observables.
@@ II.3.1 Multilayer Perceptron (MLP)
Figure 4: Using the MLP model: $p_{\rm T,jet}$, girthgirth, $M_{\rm jet}$, $N_{\rm const}$, and $Q^{\rm ch}_{\kappa}$distributions are shown for $\gamma+{\rm jet}$ with 50 << $p_{\rm T,jet}$<< 100 $\rm GeV/{\it c}$. Here, the $y$-axis is normalized by the total number of jets, NN. The blue hashed region and blue marker represent the true quark and predicted quark-like jet (from QCD Compton process); the same for red color represents gluon jet (from the annihilation process).
@@ II.3.2 Boosted Decision Tree (BDT)
In the BDT classifier, gradient boosting is employed with an ensemble of 1000 decision trees, each with a maximum depth of five. To improve stability and reduce overtraining, a learning rate (shrinkage) of 0.05 is used together with bagging, where each tree is trained on a random 80% subset of the training sample. Node splits are optimized using the Gini index as the separation criterion, and cost–complexity pruning is applied to further regularize the trees. These settings provide a balance between classification performance and robustness; the ROC is discussed in the result III.
@@ III Results and discussion
The close agreement between the BDT and MLP results indicates that the dominant discriminating information is already captured by the selected set of jet observables, with the observed performance saturation driven by physics rather than algorithmic limitations. In future, it would be important to incorporate the higher-dimensional jet substructure information to correlate QCD effects and to implement the advanced machine learning approach [20, 13, 10]. These approaches are particularly well suited to precision jet studies at the LHC and the future EIC, where controlled kinematics and reduced backgrounds enable detailed exploration of correlated QCD dynamics.